1)A hot chocolate machine adds a certain amount of chocolate into. 25 cups of hot
chocolate were investigated and found to have a mean chocolate content of 1.10
grams and a standard deviation of 0.015 grams. Find a 95% CI on the mean weight
of chocolate dispensed.(WE SHOULD USE T OR Z?WHY?)

(WE SHOULD USE T OR Z?WHY?)

It is believed that when cars go faster, their fuel consumption change. A car
company randomly assigned cars to go with 120 km/h or 90 km/h, then measured
their fuel consumption. The 180 cars that went with 120 km/h had a mean fuel
consumption of 9 liters, with a standard deviation of 0.3 liters. The 240 cars that
went with 90 km/h had a mean fuel consumption of 7 liters, with a standard
deviation of 0.4 liters. Compute a 99% confidence interval estimate for the
difference between the population means?
Answer #1

1)

We have to use t distribution because n = 25

A small sample of 25 units has a mean and a standard deviation o of 0.015. Find a 95% confidence interval of the mean Confide

Calculate high end confidence interval total: High End = X + tscorea * s/vn High End = 1.10+ 2.0639 x 0.015/125 High End = 1.

  *** Dear student we answer one question once post remaining separately****